Suppose that Mia has 10 units of labor and her two production functions are x = 6Lx and y = 10Ly. In another economy, Mr. Fantastic has 14 units of labor and production functions, x = 4Lx & y = 6Ly. In this case, the PPF for Mia is ____ and the PPF for Mr. Fantastic is ____.

Group of answer choices

4x + 6y = 280; 3x + 2y = 168

4x + 6y = 280; 2x + 5y = 180

5x + 3y = 300; 3x + 2y = 168

5x + 3y = 300; 2x + 5y = 180
Answer #1

We know Lx+Ly will be the sum of total labor

So, for Mia, 10=Lx+Ly=x/6+y/10=(5x+3y)/30 5x+3y=300

So, for Fantastic, 14=Lx+Ly=x/4+y/6=(3x+2y)/12 3x+2y=168

So, c is true.